You hold a ruler that has a charge on its tip 4 cm above a small piece of tissue paper to see if it can be picked up. The ruler has?10 �C of charge. The tissue has 1 g of mass. What is the minimum charge required to pick up the tissue paper?

(answer in �C)
Answer #1

Solution

Here in this case electrostatic force must be at least equal to the gravitational force on the piece of tissure paper.

Fe = Fg

K q1q2 / r2 = m2g

By putting the values from the given data and arranging for q2

( 9\times109)(10\times10-6) q2 = 0.001(10)

q2 = 1.1\times10-7C

q2 = 0.11\muC